how do you solve this problem in simpler like terms 0.4r-1.2=0.3r+0.6
Answer by ankor@dixie-net.com(22740)

0.4r - 1.2 = 0.3r + 0.6
:
Get rid of the decimals, multiply equation by 10
4r - 12 = 3r + 6
:
Subtract 3r from both sides:
4r - 3r - 12 = 6
:
Add 12 to both sides
r = 6 + 12
:
r = 18
:
:
Check solution in original equation, replace r with 18:
0.4(18) - 1.2 = 0.3(18) + 0.6
7.2 - 1.2 = 5.4 + .6
6 = 6; confirms our solution of r=18